A 2.0?C charge and a 2.2?C charge are 10.0cm apart. Find a point where the electric field is zero. Enter the distance from the required point to the charge 2.0?C .
Answer #1


let EF be sero at x units from 2 uC charge

so at this point as EF = 0

E 1 = E2

so

Kq1/x^2 = Kq2/(0.1-x)^2

(0.1-x)/x = sqrt(2.2/2)

0.1-x = 1.048 x

2.048 x = 0.1

x = 0.1/2.048

x = 4.88 cm      from   2uC charge , EF is zero
Answer #2

I have solved this question earlier with different figures. Please workout using yours figures. If you need any further help just PM me. If I have helped you please rate me 5 stars first (before you rate anyone else)

A 1.1 uC charge and a 3.5 uC charge are 15.0 cm apart.

Find a point where the electric field is zero.

Express your answer using two significant figures. Enter the distance from the required point to the charge 1.1 uC .

Ans

1.1/x^2 =3.5/(15-x)^2

15/x -1 =1.78

x =5.388 cms from 1.1 uC